OpenStudy (anonymous):

convert y-5=-2(X+1)^2 to quadratic intercept form

OpenStudy (jhannybean):

quadratic formula follows \(\large y= ax^2+bx+c\) and you have \(y-5=-2(x+1)^2\) first we need o expand (x+1)^(2) = (x+1)(x+1) = \(x^2+2x+1\) then multiply in the 2. \[2(x^2+2x+1)=(2x^2+4x+2)\] in order to fit the form now, we have to add 5 to both sides, so we get \[y=(2x^2+4x+2)+5\] we can simplify it further \[y=2x^2+4x+7\]All together, quadratic intercept form: \(\large y=ax^2+bx+c\) => \(\large y=2x^2+4x+7\)

OpenStudy (anonymous):

Sorry but thats standard form. I'm asking for the intercept form which is y=a(x-r)(x-r)

OpenStudy (anonymous):

y - 5 = -2 X2

OpenStudy (jhannybean):

\(\large y=2x^2+4x+7\)= standard form\[\large \frac{y}{2}=x^2+2x+7\]complete the square\[\large \frac{y}{2}=x^2+2x+1=-\frac{7}{2}+1\]\[\large \frac{y}{2}=(x+1)^2+\frac{5}{2}\]multiply by 2 the whole thing. \[\large y=[(x+1)^2+\frac{5}{2}]*2\]\[\large y=2(x+1)^2+5\] rearrange. \(\large y-5=2(x+1)^2\) intercept form.